# Semantic Web: Past, Present, and Future

### v1

@@ 8 Trustworthiness and Provenance of Graph Data
Trustworthiness of web pages and data on the web can be detected by various indicators, e. g., by certificates, by the placement of search engine results, and by links (forward and backward links) to other pages. However, on the Semantic Web, there are few ways for users to assess the trustworthiness of individual data. Rules can be utilized to define policies and business logic over the web of data, and transparently used to infer data that validate or do not validate these policies. The trustworthiness of inferred data can be assessed through its provenance, which encompasses metadata detailing how the data was acquired and verified [102].
@@
The trustworthiness of data on the web can be inferred from the trustworthiness of other users (“Who said that?”), the temporal validity of facts (“When was a fact described?”), or in terms of uncertainty of statements (“To what degree is the statement true?”). Artz and Gil [10] summarize trustworthiness as follows: “Trust is not a new research topic in computer science, spanning areas as diverse as security and access control in computer networks, reliability in distributed systems, game theory and agent systems, and policies for decision-making under uncertainty. The concept of trust in these different communities varies in how it is represented, computed, and used.” Although trustworthiness has long been considered in these areas, the provision and publication of data by many users to multiple sources on the Semantic Web introduces new and unique challenges.
@@
One way of facilitating trust on the Semantic Web is to capture and provide the provenance of data with the PROV ontology (PROV-O) [^60] https://www.w3.org/TR/prov-o/. It captures information about which Agents cause data Entities to be processed by which Activities. Capturing such information requires the use of known tools for modeling metadata for RDF data, e. g., reification, singleton properties, named graphs, or RDF-star [^61] https://w3c.github.io/rdf-star/. While some approaches use these constructs to capture provenance information for each triple individually [59], others exploit the fact that typically multiple triples share the same provenance [77] so that they can be combined into the same named graph encoding the provenance information only once for a set of triples. Delva et al. [44] introduce the notion of shape fragments, which entail the validation of a given shape through the neighborhood of a node, along with the node’s provenance and the rationale behind its validation.
@@
Furthermore, trustworthiness also plays a role in inference services on the Semantic Web, as data inference must consider specifications related to trustworthiness and data must be evaluated for trustworthiness. Important aspects for trustworthiness of data include [10]: the origin of the data, trust already gained based on previous interactions, ratings assigned by policies of a system, and access controls and, in some cases, security and importance of information. These aspects are realized in different systems.
@@
In general, data provenance and trustworthiness of data on the Semantic Web have been addressed for RDF data [47, 57] as well as for OWL and rules in [47]. In addition, there are some recent approaches on supporting how-provenance for SPARQL queries [65, 84, 58] with the goal of providing users with explanations on how the answers to their queries were derived from the underlying graphs. Other work deals with access controls over distributed data on the Semantic Web [63]. Furthermore, there are approaches to computing trust values [150] and informativeness of subgraphs [98]. There are also digital signatures for graphs [19]. Analogous to digital signatures for documents, entire graphs or selected vertices and edges of a graph are provided with a digital signature to ensure the authenticity of the data and thus detect unauthorized modifications [99]. In the approach for digital graph signatures developed by Kasten et al., graph data on the Web is supported in RDF format as well as in OWL [100]. The digital graph signature is itself represented as a graph again and can thus be published together with the data on the Web. The link between the signature graph and the signed graph is established by the named graph mechanism [100], although other mechanisms are also possible. Through this mechanism, it is possible to combine and nest signed graphs. It is thus possible to re-sign already signed graphs together with other, new graph data, etc. This makes it possible to build complex chains of trust between publishers of graph data and to be able to prove the origin of data [100, 99].
@@ 9.2 Semantic Search
Sig.ma [158] was an application for (browsing) Semantic Web data, which may come from multiple distributed data sources. Sig.ma provided an API for automatically integrating multiple data sources on the Web. The requested data sources describe information in RDF. A search in Sig.ma was initiated by a textual query from the user. Entities such as people, places, or products can be searched for. Results of a query are presented in aggregated form, that is, properties of the searched entity, such as a person, are presented in aggregated form from different data sources. For example, in a person search, information such as e-mail address, address, or current employer can be displayed. In addition to the actual information, links to the underlying data sources are also displayed to allow users to navigate to refine their search. Sig.ma also supported structured queries in which specific characteristics can be requested for an entity, such as contact information for a specific person.
Queries to data sources occur in parallel. The results from each data source in the form of RDF graphs are summarized by using properties of links in RDF data, such as owl:sameAs, or inverse-functional predicates. When searching data sources, techniques such as indexes, logical inference, and heuristics are used for data aggregation. OntoBroker [^63] https://www.semafora-systems.com/ontobroker-and-ontostudio-x [43] and OntoEdit [153] are ontology editors with search and inference systems for ontologies. Using OntoBroker, complex queries over distributed Semantic Web resources, e. g., represented in OWL, RDF, RDFS, SPARQL, and also F-Logic) can be efficiently processed.
@@ 9.3 Knowledge Graphs and Wikidata
There is an increasing number of knowledge bases and representations of structured data. For example, the secondary database Wikidata [^64] https://www.wikidata.org/wiki/Wikidata:Introduction/de [165]. A secondary database includes, in addition to the (actual) statements, relationships to their sources and other databases (called secondary information). Wikidata is a shared database between Wikipedia and Wikimedia. Wikidata mainly contains a collection of objects, which are represented as triples over the objects’ properties and the corresponding values. Semantic MediaWiki [^65] https://www.semantic-mediawiki.org/wiki/Semantic_MediaWiki is an extension of MediaWiki. It serves as a flexible knowledge base and knowledge management system. Semantic MediaWiki extends a classic wiki with the ability to enrich content in a machine-readable way using semantic annotations.
Another knowledge base was Freebase [^66] https://www.freebase.com, also an open and collaborative platform initiated in 2007 and acquired by Google in 2010. The content from Freebase was taken from various sources, including parts from the MusicBrainz ontology mentioned earlier. The success and widespread use of Wikidata prompted Google to migrate Freebase to Wikidata [154]. This strengthened the goal to develop a comprehensive, collaborative basis of structured data.

### v2

@@ 8 Trustworthiness and Provenance of Graph Data
Trustworthiness of web pages and data on the web can be detected by various indicators, e. g., by certificates, by the placement of search engine results, and by links (forward and backward links) to other pages. However, on the Semantic Web, there are few ways for users to assess the trustworthiness of individual data. Rules can be utilized to define policies and business logic over the web of data, and transparently used to infer data that validate or do not validate these policies. The trustworthiness of inferred data can be assessed through its provenance, which encompasses metadata detailing how the data was acquired and verified [DBLP:conf/rr/KhandelwalJK11].
@@
The trustworthiness of data on the web can be inferred from the trustworthiness of other users (“Who said that?”), the temporal validity of facts (“When was a fact described?”), or in terms of uncertainty of statements (“To what degree is the statement true?”). Artz and Gil [DBLP:journals/ws/ArtzG07] summarize trustworthiness as follows: “Trust is not a new research topic in computer science, spanning areas as diverse as security and access control in computer networks, reliability in distributed systems, game theory and agent systems, and policies for decision-making under uncertainty. The concept of trust in these different communities varies in how it is represented, computed, and used.” Although trustworthiness has long been considered in these areas, the provision and publication of data by many users to multiple sources on the Semantic Web introduces new and unique challenges.
@@
One way of facilitating trust on the Semantic Web is to capture and provide the provenance of data with the PROV ontology (PROV-O) [^60] https://www.w3.org/TR/prov-o/. It captures information about which Agents cause data Entities to be processed by which Activities. Capturing such information requires the use of known tools for modeling metadata for RDF data, e. g., reification, singleton properties, named graphs, or RDF-star [^61] https://w3c.github.io/rdf-star/. While some approaches use these constructs to capture provenance information for each triple individually [DBLP:conf/semweb/GalarragaMH17], others exploit the fact that typically multiple triples share the same provenance [DBLP:conf/semweb/HansenLGLTH20] so that they can be combined into the same named graph encoding the provenance information only once for a set of triples. Delva et al. [DBLP:conf/edbt/DelvaDJB23] introduce the notion of shape fragments, which entail the validation of a given shape through the neighborhood of a node, along with the node’s provenance and the rationale behind its validation.
@@
Furthermore, trustworthiness also plays a role in inference services on the Semantic Web, as data inference must consider specifications related to trustworthiness and data must be evaluated for trustworthiness. Important aspects for trustworthiness of data include [DBLP:journals/ws/ArtzG07]: the origin of the data, trust already gained based on previous interactions, ratings assigned by policies of a system, and access controls and, in some cases, security and importance of information. These aspects are realized in different systems.
@@
In general, data provenance and trustworthiness of data on the Semantic Web have been addressed for RDF data [DBLP:journals/ki/DividinoSSS09, DBLP:conf/semweb/FlourisFPTC09] as well as for OWL and rules in [DBLP:journals/ki/DividinoSSS09]. In addition, there are some recent approaches on supporting how-provenance for SPARQL queries [DBLP:journals/jacm/GeertsUKFC16, DBLP:journals/pvldb/HernandezGH21, DBLP:conf/www/Galarraga0KH23] with the goal of providing users with explanations on how the answers to their queries were derived from the underlying graphs. Other work deals with access controls over distributed data on the Semantic Web [DBLP:conf/esws/GavriloaieNOSW04]. Furthermore, there are approaches to computing trust values [DBLP:journals/jair/StoilosSPTH07] and informativeness of subgraphs [DBLP:conf/cikm/KasneciEW09]. There are also digital signatures for graphs [DBLP:conf/ecai/BellomariniNS20]. Analogous to digital signatures for documents, entire graphs or selected vertices and edges of a graph are provided with a digital signature to ensure the authenticity of the data and thus detect unauthorized modifications [DBLP:phd/dnb/Kasten16]. In the approach for digital graph signatures developed by Kasten et al., graph data on the Web is supported in RDF format as well as in OWL [DBLP:conf/esws/KastenSS14]. The digital graph signature is itself represented as a graph again and can thus be published together with the data on the Web. The link between the signature graph and the signed graph is established by the named graph mechanism [DBLP:conf/esws/KastenSS14], although other mechanisms are also possible. Through this mechanism, it is possible to combine and nest signed graphs. It is thus possible to re-sign already signed graphs together with other, new graph data, etc. This makes it possible to build complex chains of trust between publishers of graph data and to be able to prove the origin of data [DBLP:conf/esws/KastenSS14, DBLP:phd/dnb/Kasten16].
9 Machine Learning on Knowledge Graphs
Machine learning on graphs is one of the most active fields of research in AI [hamilton-book]. We will first consider graph embeddings for shallow representations of nodes, edges, and graphs. Then we turn our attention to neural networks for graphs (GNNs), which can compute more complex representations. Finally, we discuss developments in language models and their relationship to and integration with knowledge graphs.
9.1 Graph Embeddings
Graph embeddings are used to compute representations of nodes, edges, and graphs for use in machine learning tasks, such as input for GNNs. The basic idea behind embeddings is to use unsupervised learning to determine a representation of nodes, for example, via their neighborhood relationships in a high-dimensional space.
9.1.1 Classical Graph Embeddings
Classic approaches to the unsupervised calculation of graph embeddings include Node2Vec [node2vec], which uses random graph traversal to calculate how similar the start node is to nodes in its neighborhood. However, graph embeddings also include self-supervised aspects, such as in Deep Walk [deepwalk], a generalization of Node2Vec using a weighted breadth-first versus depth-first search. In Deep Walk, random traversals in the graph are also calculated starting from a node. The embeddings are then calculated using, among other things, the skip-gram idea from the embedding model Word2Vec [word2vec] for texts. The latter provides for an element to be removed from a sequence in a self-supervised approach and predicted by a neural network. In the context of Deep Walk, the sequences are the random traversals in the graph and the elements are the nodes in the graph. An overview of classical graph embeddings can be found in Hamilton [hamilton-book] and Murphy [murphy-book], among others.
9.1.2 Embeddings for Knowledge Graphs
Specific embedding models for knowledge graphs include RDF2Vec [rdf2vec] and a variety of approaches based on it. In general, embeddings for knowledge graphs can be classified based on Biswas et al. [biswas_et_al:TGDK.1.1.4] into translation-based models based on a distance function between nodes, models that model probabilities and uncertainty, and approaches based on semantic similarity determined by tensor decomposition.
Examples include TransE [transe] as a translation-based model for representing nodes and edges in a common embedding space. TransR [transr] extends TransE to support 1:N and other relationships. It uses separate embedding spaces, i. e., a separate matrix for each relation [biswas_et_al:TGDK.1.1.4], and therefore requires more memory. While TransR is limited to the immediate neighborhood, PTransR extends this to include the encoding of paths. TransG [transg] is based on a mixed model of Gaussian distributions to map different meanings of a relation such as hasPart.
While these embedding models operate on knowledge graphs, they do not map the specific semantics of Semantic Web standards such as RDFS and OWL. One approach in this direction is RotateE [rotate], which uses rotational transformations to map complex relationships in knowledge graphs [biswas_et_al:TGDK.1.1.4], such as symmetric and antisymmetric relations, inversion, and composition. Graph embedding models that can map parts of the semantics of OWL ontologies are TransOWL [transowl] or OWL2Vec [owl2vec]. These map the semantics of description logic expressions such as the intersection of two classes, subclasses, etc. directly in the training of the embeddings using special loss functions. TransOWL extends previous embedding models by encoding not only the directly observed graph, but also knowledge derived from conclusions based on available axioms. To do this, TransOWL integrates the OWL axioms directly into the loss function, effectively performing inferences during training. TransOWL is derived from TransE, while its variant TransROWL is based on TransR. In addition, there is also a version that extends TransOWL with equivalence axioms and inverse relations. Another model for OWL is OWL2Vec*, which uses random walks and word embeddings to encode the semantics of OWL ontologies.
9.1.3 Discussion of Graph Embeddings
Graph embeddings have a number of disadvantages [hamilton-book]. First, the calculations are slow due to high redundancy. Second, the embeddings of a graph cannot be easily transferred to new nodes or even new graphs without further learning. Therefore, graph embeddings are inherently transductive, meaning they assume that the entire graph is available at training time and can be used for unsupervised (Node2Vec) or self-supervised learning (Deep Walk). Therefore, the embeddings that have already been calculated cannot be transferred to new nodes. The inefficiency caused by redundant calculations stems from the fact that a separate embedding must be determined for each node via random traversals. In doing so, the same nodes are often visited multiple times for different embeddings. It should also be mentioned that the primary goal of graph embeddings is to determine a suitable representation of graphs when dataset-specific features are not available or are to be enriched. Therefore, embeddings do not usually have a specific downstream task such as node classification.
9.2 Neural Networks for Graphs
Graph Neural Networks (GNNs) extend the concept of graph embeddings by enabling more complex representations to be calculated based on the node features. In contrast to graph embeddings, GNNs are typically characterized by supervised or semi-supervised learning. The origins of GNNs can be traced back at least to the work of Gori et al. [gorietal2005] and Scarselli et al. [scarsellietal2009].
9.2.1 Categories of Graph Neural Networks (GNNs)
GNNs can be classified as isotropic and anisotropic, as well as transductive and inductive [DBLP:journals/nn/GalkeVFZHS23]. The majority of GNNs are based on the principle of message passing, in which features of neighboring nodes are used to compute representations of the nodes. In isotropic GNNs, all edges have the same influence on the calculation of node representations, while in anisotropic models, the edge weights and thus the representations of neighboring nodes are different or are learned.
As described above, transductive GNNs assume that the unlabeled nodes from the test set are also available at training time. These can be used in various calculations, for example for message passing, but have no influence on the error function of the model and are therefore not taken into account in the learning process using backpropagation. Examples of GNNs are GCN (isotropic and transductive), which is motivated by spectral analysis in graph theory [gcn]. GAT (anisotropic and inductive) learns the edge weights using attention [gat]. GraphSAGE (isotropic and inductive) uses a sampler, among other things, and thus scales for large graphs. When the hyperparameters are optimized, these models are powerful representatives of GNNs with very good performance [pitfalls, DBLP:journals/corr/abs-2406-08993-log2024].
In classical GNNs such as GCN, message passing is used to efficiently calculate embeddings for all nodes in the graph simultaneously [hamilton-book]. As mentioned above, GNNs are typically computed through supervised or semi-supervised learning and involve a downstream task such as node or graph classification, graph regression, and edge prediction (further applications are described in Murphy [murphy-book]). Learning using GNNs is often referred to as semi-supervised because in many datasets and methods, only a few nodes per class are labeled with a gold standard label [hamilton-book]; for example, less than $5\%$ of the nodes have a label when training a GCN [gcn].
9.2.2 Oversquashing and Oversmoothing in GNNs
One problem with classical GNNs is that long paths cannot be mapped, leading to oversquashing and oversmoothing [hamilton-book]. Oversquashing refers to when the representations of distant nodes are “squashed” together over only a few edges, leading to a loss of important information for calculating the representation of the target node. Oversmoothing refers to the incremental alignment of all representations in the GNN when, after many message-passing steps, the representations of the nodes lose their individual features. These problems can be effectively addressed by inserting additional learnable weights between the GNN layers, as in the GCNII model [gcnii]. Other approaches, such as Graph-MLP, no longer depend on message passing and apply contrastive learning [graph-mlp].
9.2.3 Homophilic and Heterophilic Graphs
Another discussion surrounding GNNs concerns the distinction between homophilic and heterophilic graphs, i. e., whether neighboring nodes are highly likely to belong to the same class or to different classes [DBLP:conf/iclr/0001LST22-homophily-necessary]. While models such as GCN assume that homophily is important for good performance, later work concludes that graphs can also be heterophilic as long as the neighborhood distribution of classes is consistent [DBLP:conf/iclr/0001LST22-homophily-necessary]. A good overview of the most important GNNs can be found in Hamilton [hamilton-book] and Murphy [murphy-book], among others.
9.2.4 GNNs for Knowledge Graphs
Specific neural networks for knowledge graphs include R-GCN [r-gcn], an extension of GCN for learning relation-specific adjacency matrices. Since the number of relations in knowledge graphs can be very large [DBLP:journals/dpd/GottronKS15], R-GCN approximates the relation-specific adjacency matrices using a linear combination of fewer base matrices. KGAT is a graph attention neural network based on a hybrid combination of user-item graphs and knowledge graphs [kgat]. The goal is to model higher-order relationships such as user-book-author-user and user-book-genre-user, and it is used for recommendation tasks.
A detailed overview of neural networks for knowledge graphs can be found, e. g., in Kumar et al. [DBLP:journals/access/YeKSSW22-kg-survey]. A discussion of future topics can be found in d’Amato et al. [DBLP:journals/tgdk/dAmatoMMS23].
10 Language Models and Knowledge Graphs
Language models based on neural networks, in particular the Transformer architecture [DBLP:conf/nips/VaswaniSPUJGKP17], are often distinguished into encoder models (e. g., BERT [bert]), encoder-decoder models (such as T5 [t5]), and decoder models (e. g., Llama [llama3] and GPT [gpt4]). These differ in their capabilities and primary uses. Encoder models are particularly suitable for classification [DBLP:conf/acl/GalkeS22], while decoder models dominate the field of generative AI and chatbots [are-we-makin-progress]. However, all models share characteristics that are summarized under the term foundation models [foundation-models]. These include pre-training using self-supervised learning, the use of a broad corpus as training data, and the associated applicability or adaptability to a variety of downstream tasks. The size of the language models ranges from several hundred million parameters (BERT) to billions of parameters (Llama, GPT). These are therefore also referred to as small language models versus large language models.
10.1 Graph Transformer, Language Models, and Retrieval Augmented Generation
Adaptations of the transformer architecture to graphs, such as graph transformer networks [graph-transformer-networks], have already shown strong performance on downstream tasks in 2021. However, initial approaches were not yet based on pre-training models and were therefore not transferable to different downstream tasks. One reason for this is that the features between the datasets are too specific and also have different dimensionalities. Some early approaches to integrating knowledge graphs with language models have therefore aimed to encode the facts from knowledge graphs directly into the parameters of language models. Interesting approaches in this direction include SKILL [DBLP:conf/naacl/MoiseevDAJ22-skill], in which the language model T5 was trained on a textual representation of factual knowledge from knowledge graphs. SKILL shows that pre-training T5 on the Wikidata knowledge graph [DBLP:journals/cacm/VrandecicK14] leads to better downstream performance. The knowledge graph “disappears” into the model parameters. Changing the knowledge graph or deleting facts therefore becomes very laborous. ATLAS [DBLP:journals/jmlr/IzacardLLHPSDJRG23-atlas] is a retrieval-augmented method (RAG) for integrating knowledge-intensive tasks into a language model with few training examples. The goal of ATLAS is to learn facts and use them, e. g., for fact checking or asking trivia questions. An overview of methods for editing knowledge graphs in language models can be found, e. g., in Cao et al. [DBLP:journals/ijautcomp/CaoLHS24].
Due to the text-based nature of language models such as T5, it was necessary for the graphs to be represented in text form. Features such as TF-IDF in GCN datasets cannot be well represented in so-called text-attributed graphs. Knowledge graphs in RDFS or OWL, on the other hand, are well suited due to their descriptive properties of nodes and edges. A collection of text-based graphs can be found in TAGLAS [DBLP:journals/corr/abs-2406-14683-taglas], among others.
10.2 Graph Foundation Models
Under the term graph foundation models, there are a number of approaches for graphs that meet the above requirements. The models differ in terms of their architecture, i. e., whether they are based on an existing language model, a GNN, or a combination of language model and GNN [graph-foundation-models-survey]. The Graph Language Model (GLM) [glm], which is based on Google’s T5 language model, belongs to the first category. The attention mechanism of T5’s transformer layer has been modified so that masking takes into account the adjacency of nodes. This modification of the attention mask also distinguishes GLM from the above-mentioned approaches such as SKILL. For T5, the input graph is converted into a Levi graph and then processed as text, like any other prompt, in the T5 model. Due to the nature of GLM, it is suitable for processing both text and graphs natively, but is limited to text-attributed graphs.
Models such as PRODIGY, PSP, and GraphAny pursue an approach based on GNNs. The first two, PRODIGY [prodigy] and PSP [psp], can be understood as native graph-based foundation models that map a graph prompt in analogy to a text prompt. Based on a pre-trained model, similar to language models, several graphs can be given as examples (context learning) along with the task to be solved. Unlike PRODIGY, PSP works with prototypes that represent the graphs. GraphAny [graphany] uses a combination of different linear GNNs based on the Simplified GCN Model (SGC) [sgc]. In order to meet the requirement of being usable for various downstream tasks, GraphAny has a mechanism for handling different dimensionalities of node features in graph datasets. GraphAny is very efficient due to the use of linear models. In addition, training on a few examples is sufficient to achieve strong performance on other datasets with different features. Approaches in the field of combining language models and GNNs include [opengraph, ofa, gofa].
An overview of graph foundation models is provided by Liu et al. [DBLP:journals/corr/abs-2310-11829-graphfoundationmodels-survey] and Fan et al. [DBLP:journals/corr/abs-2404-14928].
@@ 11.2 Semantic Search
Sig.ma [key:sigma] was an application for (browsing) Semantic Web data, which may come from multiple distributed data sources. Sig.ma provided an API for automatically integrating multiple data sources on the Web. The requested data sources describe information in RDF. A search in Sig.ma was initiated by a textual query from the user. Entities such as people, places, or products can be searched for. Results of a query are presented in aggregated form, that is, properties of the searched entity, such as a person, are presented in aggregated form from different data sources. For example, in a person search, information such as e-mail address, address, or current employer can be displayed. In addition to the actual information, links to the underlying data sources are also displayed to allow users to navigate to refine their search. Sig.ma also supported structured queries in which specific characteristics can be requested for an entity, such as contact information for a specific person.
Queries to data sources occur in parallel. The results from each data source in the form of RDF graphs are summarized by using properties of links in RDF data, such as owl:sameAs, or inverse-functional predicates. When searching data sources, techniques such as indexes, logical inference, and heuristics are used for data aggregation. OntoBroker [^63] https://www.semafora-systems.com/ontobroker-and-ontostudio-x [Decker1999] and OntoEdit [10.1007/3-540-48005-6_18] are ontology editors with search and inference systems for ontologies. Using OntoBroker, complex queries over distributed Semantic Web resources, e. g., represented in OWL, RDF, RDFS, SPARQL, and also F-Logic) can be efficiently processed.
@@ 11.3 Knowledge Graphs and Wikidata
There is an increasing number of knowledge bases and representations of structured data. For example, the secondary database Wikidata [^64] https://www.wikidata.org/wiki/Wikidata:Introduction/de [DBLP:journals/cacm/VrandecicK14]. A secondary database includes, in addition to the (actual) statements, relationships to their sources and other databases (called secondary information). Wikidata is a shared database between Wikipedia and Wikimedia. Wikidata mainly contains a collection of objects, which are represented as triples over the objects’ properties and the corresponding values. Semantic MediaWiki [^65] https://www.semantic-mediawiki.org/wiki/Semantic_MediaWiki is an extension of MediaWiki. It serves as a flexible knowledge base and knowledge management system. Semantic MediaWiki extends a classic wiki with the ability to enrich content in a machine-readable way using semantic annotations.
Another knowledge base was Freebase [DBLP:conf/sigmod/BollackerEPST08-freebase, DBLP:conf/aaai/BollackerCT07-freebase], also an open and collaborative platform initiated in 2007 and acquired by Google in 2010. The content from Freebase was taken from various sources, including parts from the MusicBrainz ontology mentioned earlier. The success and widespread use of Wikidata prompted Google to migrate Freebase to Wikidata [DBLP:conf/www/TanonVSSP16]. This strengthened the goal to develop a comprehensive, collaborative basis of structured data.